OpenStudy (anonymous):

using Demoivre's theorem solve this equation .. (2+2i)^6

OpenStudy (amistre64):

thats the z(cos sin ) stuff right?

OpenStudy (amistre64):

-512i if your lucky :)

OpenStudy (amistre64):

using ( ) as delimiters instead of [ ] places it inline

OpenStudy (amistre64):

if we type it out like this \(\frac{a}{b}\) it should go inline text

OpenStudy (anonymous):

thank you but how can i work it out?

OpenStudy (2bornot2b):

work what out?

OpenStudy (anonymous):

i dont understand... doesnt the 2^6 have to be a square root?

OpenStudy (amistre64):

2^6 = 1 2 4 8 16 32 64.... 64 cos(6pi) = 1 and sin(6pi)= 0 how the wolf gets -516i from that i cant tell yet

OpenStudy (amistre64):

isnt cos(2pi) = 1?

OpenStudy (amistre64):

1+i = pi/4 i believe is a better item

OpenStudy (amistre64):

1 over 1i up is a 45 tri

OpenStudy (amistre64):

or i might be off in my thinking there

OpenStudy (amistre64):

\[(2+2i)^6=2^6(\cos(\pi/4)+i\sin(\pi/4))^6=2^6(\cos(6\pi/4)+i\sin(6\pi/4))\] 64(\cos(3\pi/2)+i\sin(3\pi/2))\] \[64(0-i)\] ugh

OpenStudy (amistre64):

i got the 270^o yay!!

OpenStudy (amistre64):

apparently the radius aint 2

OpenStudy (anonymous):

but how do you get -512i from 64(0-i)?

OpenStudy (amistre64):

the 2^6 aint correct; the (0-i) parts good

OpenStudy (amistre64):

or at least thats what im having an issue with

OpenStudy (amistre64):

2^9 = 512

OpenStudy (amistre64):

that still gives a radius of 64 and not 512 tho

OpenStudy (earthcitizen):

\[(\sqrt{8})^6 <(45*6 )\]

OpenStudy (anonymous):

where do you get sqrt 8?

OpenStudy (amistre64):

the original radius is 2sqrt(2); not 2 [2sqrt(2)]^6 = 64*8 = 512

OpenStudy (amistre64):

we have to power up the origonal radius

OpenStudy (amistre64):

i knew if i throew demovers out the window I could think of it lol

OpenStudy (amistre64):

\[r^n(cos(nx)+isin(nx))\]

OpenStudy (earthcitizen):

\[r ^{n}<n \theta\]

OpenStudy (2bornot2b):

\[(2+2i)=2\sqrt2(\cos(\pi/4)+i\sin(\pi/4))\] \[(2\sqrt2)^6(\cos6(\pi/4)+i\sin6(\pi/4))\]

OpenStudy (amistre64):

tada!! :) i knew you could do it class is starting so ciao

OpenStudy (anonymous):

thanks! i kinda get it now.. but where does the 6 come from after the cos and sin

OpenStudy (earthcitizen):

\[r=\sqrt{2^{2}+2^{2}}=\sqrt{4+4}=\sqrt{8}\] arg z =arctan(1)=45 \[(\sqrt{8}<45)^{6}\]

OpenStudy (anonymous):

oh wow alright thank you

OpenStudy (earthcitizen):

\[r ^{n}<n \theta \] r=(8^1/2)^6=512, n=6 and theta=45*6=270 512<270\[r(\cos \theta+isin \theta) = 512(\cos(270)+isin(270))=-i512\]